Cars are equipped with flexble bumpers, so that they will bounce off each other during low-speed collisions. In one particular accident, an 1800 kg car traveling at 1.60 m/s to the left collides with a 1300 kg car that is moving at 1.25 m/s to the right. If the lighter car moves at 0.520 m/s to the left just after the collision and ignoring any road friction, (a) what is the speed of the heavier car just after the collision? (b) What is the change in the combined kinetic energy of the two-car system during this collision?

1
Expert's answer
2022-04-22T13:53:44-0400

"m_1v_1+m_2v_2=m_1u_1+m_2u_2,"

"u_2=\frac{m_1(v_1-u_1)-m_2v_2}{m_2}=2.5~\frac ms,"

"\u2206E=E_2-E_1=0."